what is the midpoint of the segment joint (3,5) and (-6,11)?
Answer by Alan3354(69443)

what is the midpoint of the segment joint (3,5) and (-6,11)?
--------------------
Find the average of x and of y
x: (3-6)/2 = -3/2
y: (5+11)/2 = 8
--> (-3/2,8) midpoint